### v1

@@ Appendix A Appendix
The submodularity of the performance function gives rise to the countermonotone coupling of proposition 3. The following results are proven, among others, by RachevRueschendorf1998.
@@
First, note that simple algebra, combined with assumption 2, yields $|B|<1$, and hence $1-B^{2}>0$. Hence, it is sufficient to prove that $N\equiv\tilde{R}[1-(B)^{2}]-(A^{1}+BA^{2})\geq 0$ and $M\equiv\tilde{R}[1-(B)^{2}]-(A^{2}+BA^{1})\geq 0$. We start with the first inequality. The expression $N$ is a polynomial in $(p,q,\tilde{R},\tilde{r},\beta)$. Consider the closed parameter set defined by
$p\geq 0,\quad 1-p\geq 0,$
$q\geq 0,\quad p-q\geq 0,$
$\tilde{r}\geq 0,\quad\tilde{R}-\tilde{r}\geq 0,$
$\beta\geq 0,\quad 1-\beta\geq 0,$
$\tfrac{1}{4}-\tilde{R}\geq 0,$
where the last constraint is simply a restatement of assumption 2 in our context. Using a sum-of-squares (SOS) relaxation [^14] The relaxation is computed with SumOfSquares.jl and the SCS solver. of degree $6$ we obtain an explicit identity
$4N=\nu_{0}+\sum_{i=1}^{9}\nu_{i}g_{i},$
where $g_{i}$ are the constraint polynomials above and each $\nu_{i}$ is a sum of squares. Since $\nu_{i}\geq 0$ and $g_{i}\geq 0$ on the feasible set, the right-hand side is nonnegative, implying $N\geq 0$. This proves that $N\geq 0$. $M\geq 0$ is proved in the same way. The expression $M$ is a polynomial in $(p,q,\tilde{R},\tilde{r},\beta)$, defined on the same closed parameter set reported above. An SOS relaxation of degree $6$ returns an explicit certificate
$4M=\mu_{0}+\sum_{i=1}^{9}\mu_{i}g_{i},$
with each $\mu_{i}$ a sum-of-squares polynomial. Therefore $M\geq 0$ on the feasible set. Remark 1 implies that, if an equilibrium where both players work on different technology exists, the maximal level of aggregate first-period effort exerted in the first period is $e_{1}(1,1)=2\tilde{R}$.
@@
Summarizing our findings starting from this final step of the proof, we have shown that, when we restrict attention to beliefs in $L$, a team with initial beliefs $p^{O}$ maximizes the equilibrium first-period aggregate effort, and therefore the expected first-period output evaluated at belief $p^{\star}=(x,x)$. In addition, we have shown that belief profile $p^{O}$ admits a most productive equilibrium where players invest this maximal effort level on the same technology, maximizing the expected second period output when evaluated at $t=1$ according to belief profile $p^{\star}$. Finally, we have shown that $p^{O}$ is the unique maximizer of such expected second-period output if $x<1$. This completes the proof of lemma A.1. $\blacksquare$

### v2

@@ Appendix A Appendix
The submodularity of the performance function gives rise to the countermonotone coupling of proposition 3. The following results are proven, among others, by Rachev and Rüschendorf (1998).
@@
The remark follows from the comparative statics results by Milgrom and Roberts (1990). To apply those results, we focus on the reduced-form effort-choice game played in period $t=1$ when player $1$ works on technology $A$ and player $2$ works on technology $B$. Let $\eta=-e^{2}_{1}$ and, for each $i$, define $v_{i}(e^{1}_{1},\eta;p,q)$ to be $i$’s expectation of her total payoff in the game when the team has a common prior $(p,q)$, first-period effort levels are given by $(e^{1}_{1},\eta)$ and second-period effort levels follow equilibrium play. Note that $B=-\frac{3}{2}\beta(1-p)q\left[\Delta^{2}(p+q)+2r\Delta\right]\in(-1,0]$. Hence $\frac{\partial^{2}v_{i}(e^{1}_{1},\eta;p,q)}{\partial e^{1}_{1}\partial\eta}=-B\geq 0$ for $i=1,2$, and the game is supermodular. We now show that $\frac{\partial^{2}v_{1}(e^{1}_{1},\eta;p,q)}{\partial e^{1}_{1}\partial p}\geq 0$ and $\frac{\partial^{2}v_{2}(e^{1}_{1},\eta;p,q)}{\partial\eta\partial p}\geq 0$. Note that $\frac{\partial^{2}v_{1}(e^{1}_{1},\eta;p,q)}{\partial e^{1}_{1}\partial p}=\frac{\partial A^{1}}{\partial p}+\gamma\frac{\partial B}{\partial p}$ for some $\gamma\in[0,1]$, and $\frac{\partial^{2}v_{2}(e^{1}_{1},\eta;p,q)}{\partial\eta\partial p}=-\frac{\partial A^{2}}{\partial p}-\delta\frac{\partial B}{\partial p}$, for some $\delta\in[0,1]$. Under our parameter restrictions — in particular, by assumption 2 — we have that $\frac{\partial A^{1}}{\partial p}>0$ and $\frac{\partial A^{1}}{\partial p}+\frac{\partial B}{\partial p}>0$, implying $\frac{\partial^{2}v_{1}(e^{1}_{1},\eta;p,q)}{\partial e^{1}_{1}\partial p}>0$. Additionally, we have $\frac{\partial A^{2}}{\partial p}\leq 0$ and $\frac{\partial A^{2}}{\partial p}+\frac{\partial B}{\partial p}\leq 0$, implying $\frac{\partial^{2}v_{2}(e^{1}_{1},\eta;p,q)}{\partial\eta\partial p}\geq 0$. Given that $\frac{\partial^{2}v_{1}(e^{1}_{1},\eta;p,q)}{\partial e^{1}_{1}\partial p}\geq 0$ and $\frac{\partial^{2}v_{2}(e^{1}_{1},\eta;p,q)}{\partial\eta\partial p}\geq 0$, theorem 6 by Milgrom and Roberts (1990) implies that the equilibrium level of $e^{1}_{1}(p,q)$ is non-decreasing in $p$. Hence, for every $p\in[0,1]$ and $q\in[0,p]$, we have that player $1$’s effort level satisfies $e^{1}_{1}(p,q)\leq e^{1}_{1}(1,q)=\tilde{R}$, which proves inequality 19. To prove 20 we show that if $q\leq p$ then $e^{2}_{1}(p,q)\leq e^{1}_{1}(p,q)$. Note that, using 17 and 18, we get $e^{1}_{1}(p,q)-e^{2}_{1}(p,q)\gtreqless 0$ if $A^{1}-A^{2}\gtreqless 0$. But simple algebra yields $A^{1}-A^{2}=(p-q)\Delta\left[1+\frac{3}{2}\beta(1-p)\Delta(1-q)\right]\geq 0$, where the inequality follows from $0<q\leq p$. Hence $e^{2}_{1}(p,q)\leq e^{1}_{1}(p,q)\leq\tilde{R}$, which proves inequality 20.
@@
Remark 1 implies that, if an equilibrium where both players work on different technology exists, the maximal level of aggregate first-period effort exerted in the first period is $e_{1}(1,1)=2\tilde{R}$. Summarizing our findings starting from this final step of the proof, we have shown that, when we restrict attention to beliefs in $L$, a team with initial beliefs $p^{O}$ maximizes the equilibrium first-period aggregate effort, and therefore the expected first-period output evaluated at belief $p^{\star}=(x,x)$. In addition, we have shown that belief profile $p^{O}$ admits a most productive equilibrium where players invest this maximal effort level on the same technology, maximizing the expected second period output when evaluated at $t=1$ according to belief profile $p^{\star}$. Finally, we have shown that $p^{O}$ is the unique maximizer of such expected second-period output if $x<1$. This completes the proof of lemma A.1. $\blacksquare$